Solve by completing the square.
x^2=5x+2
Please help.
Thanks.

Answer by mukhopadhyay(490)

x^2=5x+2
=> x^2-5x = 2
=> x^2 - 5x + (5/2)^2 = 2 +(5/2)^2
=> (x-5/2)^2 = 33/4
=> (x-5/2) = sqrt(33)/2 or (x-5/2) = -sqrt(33)/2
=> x = 5/2 + sqrt(33)/2 or x = 5/2 - sqrt(33)/2